x=?
in Equation x^2+4=9

Found 2 solutions by Aswathy, mathstutor494:
Answer by Aswathy(23)

x^2+4=9
x^2=9-4
x^2=5
x=+/-sqrt(5)

Answer by mathstutor494(120)

x^2+4=9
x^2 = 9-4
x^2=5
Hence x= +-sqrt(5)